# A Backdoor Attack Scheme with Invisible Triggers Based on Model Architecture Modification

### v2

@@ Abstract
Index Terms:
Backdoor Attacks, Model Architecture, Invisible Triggers, Data Manipulation, Security Risks
@@ III-A Threat model
We assume that an attacker aims to influence the training process of a neural network, where the user receives a model from the attacker. This scenario could arise when the user downloads a pre-trained model from the internet or outsources model training to a third party, both of which are very common in reality. The attacker’s goal is to generate a model with a backdoor so that it produces results significantly different from normal predictions when specific triggers are present in the input images, while concealing the existence of the backdoor and triggers. This paper describes a backdoor attack initiated solely through the model architecture, proposing a new method for covertly implanting backdoors. This method does not rely on the original model’s weights and can be directly added to a well-trained model to take effect.In this section, we use a simple example based on ResNet-18[26] to explain the design. Note that, in practice, this scheme can be injected into any architecture. The process can be divided into three stages: architecture modification, trigger generation, and backdoor activation.
@@
It is well known that the original range of values for RGB images is [0, 255]. When using deep learning frameworks like PyTorch[27], the input data is preprocessed, and transforms.ToTensor() is always used, which normalizes the original pixel values to a range of [0, 1] as floating-point numbers. Our plan is to perform simple addition or subtraction by 1 on the pixel values of the original image. The goal is that, when the user receives the model and dataset, the pixel values of the images will be normalized through preprocessing. After normalization, we multiply the pixel values by a certain factor and round them, ensuring that almost all pixel values in the image become either odd or even. As shown in Algorithm 1. We use this overall odd or even pattern as the trigger, ensuring its high stealthiness and making it completely undetectable by the naked eye or detection tools.As shown in Figure 2, the two images are indistinguishable from each other.
Input : The selected dataset $I$ contains multiple images, where each image has a shape of $(C,H,W)$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width. The dataset is chosen to include images from the class that needs backdoor injection.
@@ 18                               $I^{\prime}[i,c,h,w]\leftarrow I[i,c,h,w]$;
Data Processing. A copy of the input image is made, and all its values are multiplied by 10000 and rounded. Why choose 10000? Let the initial input data be x Through extensive experimentation, we have found that when the original pixel value x is normalized to $\frac{x}{255}$, even if rounding $\left(\frac{x}{255}\times 10000\right)$ does not result in an even number, rounding $\left(\frac{x-1}{255}\times 10000\right)$ or $\left(\frac{x+1}{255}\times 10000\right)$ will inevitably result in an even number. Multiplying by a factor less than 10000 (such as 1000 or 100) would not guarantee that adding or subtracting 1 from the original pixel value x would result in an even number. On the other hand, multiplying by a factor greater than 10000 would make the values too large and computationally inefficient. Therefore, 10000 is an appropriate magnification factor.
[…]
@@ 16      $output\leftarrow input1+input2$
20Function DataProcessing($input$)
@@ 22      $output\leftarrow\lfloor(input*10000)\rfloor$
26Function TriggerDetector($input$)
[…]
28      return $output$
@@ IV-A Effectiveness Test
CIFAR10 Poisoning One Class 73%
CIFAR10 Poisoning Two Class 64%
CIFAR10 Poisoning Three Class 56%
CIFAR10 Poisoning Four Class 49%
CIFAR10 Poisoning Five Class 41%
@@
SVHN Poisoning One Class 79%
SVHN Poisoning Two Class 69%
SVHN Poisoning Three Class 59%
SVHN Poisoning Four Class 51%
SVHN Poisoning Five Class 43%
@@
GTSRB Poisoning Five Class 75%
GTSRB Poisoning Ten Class 55%
GTSRB Poisoning Fifteen Class 40%
GTSRB Poisoning Twenty Class 33%
GTSRB Poisoning Twenty-Five Class 26%
@@ 17                         $C\leftarrow C\cup\{i\}$
CIFAR10 Poisoning One Class 83%
CIFAR10 Poisoning Two Class 74%
CIFAR10 Poisoning Three Class 53%
CIFAR10 Poisoning Four Class 57%
CIFAR10 Poisoning Five Class 49%

### v3

# Backdoor Attack with Invisible Triggers Based on Model Architecture Modification

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ II-B Backdoor Defense
Figure 1: The attacker downloads a pre-trained model from the Internet, adds a backdoor to the model, and then black-boxes it before re-uploading it to the Internet for users to download and use. Since the model has been black-boxed, users are unable to detect the presence of the backdoor architecture.
Figure 2: Users automatically combine customized models from multiple existing model architectures through Neural Architecture Search (NAS). If some of these models contain backdoor architectures, they may become part of the final model. Since a large number of models are required for NAS, it may not be possible to accurately examine the architecture of each model.
@@ III-A Threat model
We assume that an attacker aims to influence the training process of a neural network, where the user receives a model from the attacker. This situation may occur when users download pre-trained models from the Internet or outsource model training to a third party, as shown in Figure 1, or when users automatically combine customized models from multiple existing model architectures through Neural Architecture Search (NAS), as shown in Figure 2. Both scenarios are very common in reality. The attacker’s goal is to generate a model with a backdoor so that it produces results significantly different from normal predictions when specific triggers are present in the input images, while concealing the existence of the backdoor and triggers. This paper describes a backdoor attack initiated solely through the model architecture, proposing a new method for covertly implanting backdoors. This method does not rely on the original model’s weights and can be directly added to a well-trained model to take effect.In this section, we use a simple example based on ResNet-18[26] to explain the design. Note that, in practice, this scheme can be injected into any architecture. The process can be divided into three stages: architecture modification, trigger generation, and backdoor activation.
@@ III-C Trigger generation
It is well known that the original range of values for RGB images is [0, 255]. When using deep learning frameworks like PyTorch[27], the input data is preprocessed, and transforms.ToTensor() is always used, which normalizes the original pixel values to a range of [0, 1] as floating-point numbers. Our plan is to perform simple addition or subtraction by 1 on the pixel values of the original image. The goal is that, when the user receives the model and dataset, the pixel values of the images will be normalized through preprocessing. After normalization, we multiply the pixel values by a certain factor and round them, ensuring that almost all pixel values in the image become either odd or even. As shown in Algorithm 1. We use this overall odd or even pattern as the trigger, ensuring its high stealthiness and making it completely undetectable by the naked eye or detection tools.As shown in Figure 4, the two images are indistinguishable from each other.Figure 5 shows the PSNR and SSIM[28] values of 20 groups of original images and images with triggers added. It can be seen that the PSNR values are all greater than 50, and the SSIM[28] values are all greater than 0.99, indicating that there is almost no difference between the original images and the images with triggers added.
Figure 5: The image on top shows the PSNR values, while the image at the bottom shows the SSIM values.
Input : The selected dataset II contains multiple images, where each image has a shape of $(C,H,W)$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width. The dataset is chosen to include images from the class that needs backdoor injection.
@@ III-D Backdoor activation
Data Processing. A copy of the input image is made, and all its values are multiplied by 10000 and rounded. Why choose 10000? Let the initial input data be x, x is the set of all pixels in an image $\{x_{1},x_{2},x_{3},\dots,x_{n}\}$, where n is the total number of pixels in the image. Through extensive experimentation, we have found that when a particular pixel value $x_{i}$ is normalized to $\frac{x_{i}}{255}$, even if rounding $\left(\frac{x_{i}}{255}\times 10000\right)$ does not result in an even number, rounding $\left(\frac{x_{i}-1}{255}\times 10000\right)$ or $\left(\frac{x_{i}+1}{255}\times 10000\right)$ will inevitably result in an even number. Multiplying by a factor less than 10000 (such as 1000 or 100) would not guarantee that adding or subtracting 1 from the original pixel value x would result in an even number. On the other hand, multiplying by a factor greater than 10000 would make the values too large and computationally inefficient. Therefore, 10000 is an appropriate magnification factor.
$A=\frac{1}{{e}^{-\alpha\left(\left(\underset{i=0}{\overset{n}{\sum}}\frac{1+\mathrm{\cos}\left(\pi x_{i}\right)}{2}\right)-\beta\right)}}$ (1)
@@ 16 $output\leftarrow input1+input2$
20Function DataProcessing(inputinput)
@@ 22 $output\leftarrow\lfloor(input*10000)\rfloor$
26Function TriggerDetector(inputinput)
27 $output\leftarrow\frac{1}{{e}^{-\alpha\left(\left(\underset{i=0}{\overset{n}{\sum}}\frac{1+\mathrm{\cos}\left(\pi*input_{i}\right)}{2}\right)-\beta\right)}}$
28 return outputoutput
@@ IV-A Effectiveness Test
Poisoning One Class 73%
Poisoning Two Class 64%
Poisoning Three Class 56%
Poisoning Four Class 49%
Poisoning Five Class 41%
@@
Poisoning One Class 79%
Poisoning Two Class 69%
Poisoning Three Class 59%
Poisoning Four Class 51%
Poisoning Five Class 43%
@@
Poisoning One Class 75%
Poisoning Two Class 55%
Poisoning Three Class 40%
Poisoning Four Class 33%
Poisoning Five Class 26%
@@ 17          $C\leftarrow C\cup\{i\}$
Poisoning One Class 83%
Poisoning Two Class 74%
Poisoning Three Class 53%
Poisoning Four Class 57%
Poisoning Five Class 49%